OpenStudy (anonymous):

What is the distance between P(–4, 3) and Q(6, 1)? Round to the nearest tenth

OpenStudy (anonymous):

distance formula

OpenStudy (sriram):

sqrt(104z)

myininaya (myininaya):

\[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (anonymous):

10^2 + 2^2 = 104 sqrt(104)

OpenStudy (sriram):

sqrt(104)

OpenStudy (sriram):

d z wz mistake

myininaya (myininaya):

\[\text{ where} (x_1=-4,y_1=3) \text{ and} (x_2=6, y_2=1)\]

OpenStudy (anonymous):

pythagoras. from -4 to 6 is 10 from 3 to 1 is 2 \[10^2+2^2=d^2\] \[d=\sqrt{100+4}\]

OpenStudy (anonymous):

\[\sqrt{104}\] = 10.198039027185 ans : 10.2